# A Dual-Perspective NLG Meta-Evaluation Framework with Automatic Benchmark and Better Interpretability

### v1

@@ 1 Introduction
Natural Language Generation (NLG) has long been crucial in natural language processing research, encompassing common tasks such as text summarization and dialogue response generation. Unlike other tasks, such as question answering and mathematical reasoning, which have a single standard answer, making it convenient to evaluate the model output by direct matching, NLG tasks are inherently open-ended and lack a unique correct output. Therefore, they require more robust and flexible evaluation metrics. Traditional metrics primarily include string matching-based methods, such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), as well as semantic representation-based methods like BERTScore (Zhang et al., ) and BARTScore (Yuan et al., 2021). However, with the advancements of language models, these metrics have gradually been found insufficient to meet requirements. Moreover, some studies have highlighted their deficiencies in robustness.
@@
Nevertheless, we argue that traditional NLG meta-evaluation used to assess the performance of metrics has certain limitations, preventing a comprehensive understanding of their evaluation capabilities. Firstly, when measuring the consistency between metrics and human evaluations, multiple human ratings for each sample are typically averaged directly. However, human ratings are not continuous values and may be inconsistent among different annotators, and the quality intervals between adjacent ratings are not uniform (Sullivan and Artino Jr, 2013). Therefore, this seemingly intuitive averaging aggregation is not such reasonable. In addition, the choice of correlation measures commonly used in traditional meta-evaluation can significantly impact the performance assessment of different metrics (Perrella et al., 2024; Gao et al., 2024). It is unclear which correlation measure should be selected in different NLG evaluation scenarios. Furthermore, many widely used NLG evaluation benchmarks suffer from issues including outdated generation systems and potential data contamination, further undermining the reliability of meta-evaluation. More detailed discussions of these limitations are included in Section 3.
To address these issues, we propose a new dual-perspective NLG meta-evaluation framework. Specifically, the meta-evaluation from the global perspective formulates evaluation as an ordinal classification task, assessing the ability of metrics to judge targets into coarse-grained quality levels. On the other hand, the local-perspective meta-evaluation formulates evaluation as an adjacent pairwise comparison task, assessing the metrics in distinguishing between targets with fine-grained quality differences. These two perspectives focus on different evaluation capabilities, forming a more interpretable NLG meta-evaluation approach compared to the traditional meta-evaluation.
Furthermore, we introduce a method of automatic benchmark construction to better implement our proposed dual-perspective meta-evaluation. Our new benchmark can be automatically constructed based on any existing NLG evaluation benchmark without the cost of new human annotations and potential data contamination. To comprehensively assess the evaluation capabilities of emerging LLM-as-a-Judge, we further experiment with 16 representative LLMs, including both the general-purpose and specifically fine-tuned ones, based on our new NLG meta-evaluation framework. The results indicate LLMs have their respective proficiency in different evaluation capabilities, demonstrating the effectiveness and necessity of our framework.
@@
    We conduct extensive experiments with 16 representative LLMs, analyzing their respective proficiency in different evaluation capabilities. The code and data of our meta-evaluation framework will be released to facilitate future research in NLG evaluation.
@@ 3 Dual-Perspective Meta-Evaluation
Recent studies (Perrella et al., 2024; Gao et al., 2024) have investigated the impact of different consistency measures on NLG meta-evaluation, but less attention has been paid to the aggregation method. The implementation of directly averaging ratings from multiple human evaluators has been followed by default in the previous related work. However, we have doubts about this seemingly self-evident approach from two main aspects. We take the widely-used benchmark, SummEval (Fabbri et al., 2021), as an example, where human evaluation is conducted using a 5-point Likert scale with three annotators.
@@
The additional quality levels generated by averaging aggregation are not necessarily valid. The averaged ratings with the 1-5 scale and three annotators can have up to 13 distinct values, leading to some new ratings besides the original five ratings, such as rating $\frac{4}{3}$. To verify whether these new ratings really reflect finer-grained quality levels beyond the original ratings, we conduct a human re-evaluation through pairwise comparison on SummEval. Specifically, we sample specific target pairs as the first test group, where the two averaged ratings involved in each pair are adjacent and belong to new and original ratings, respectively, such as ratings $\frac{4}{3}$ and 1. As a contrast, we additionally sample the second test groups, where the two averaged ratings involved are adjacent original ratings, such as ratings 2 and 1. More experimental details are included in Appendix A.1. The proportions of pairs with consistent judgments between our re-evaluations and the averaged existing ratings are only 42% for the first group, while 88% for the second group. The results indicate that there are indeed definite quality differences among the original ratings, whereas this is not the case if adding the new ratings generated by averaging aggregation.
@@ 3.1 Global Perspective
$\kappa(\{h_{ij}\}_{i=1,j=1}^{n,m},\{x_{ij}\}_{i=1,j=1}^{n,m},l)$
where $\kappa$ represents the linear weighted $\kappa$ (Cohen, 1968), which is the most recommended ordinal classification metric by Sakai (2021) through comprehensive analyses, with details included in Appendix A.2. And $l$ represents the number of different coarse-grained ratings.
@@ 3.2 Local Perspective
On the other hand, in local-perspective meta-evaluation, the focus shifts to the ability to perceive fine-grained quality differences, thereby complementing the global perspective and rewarding those metrics with good discriminative ability. More specifically, given the targets with different qualities, where some of them may have the same coarse-grained rating, the concern here is whether they can be correctly ranked rather than whether they can be rated correctly. Furthermore, through a target sequence for each source that contains significantly more than $l$ targets with different qualities, we can enable the more fine-grained meta-evaluation. The performance of metrics is assessed by comparing their evaluation scores for each pair of adjacent targets in the sequence sorted by target quality, which is the most challenging case. Given a target sequence $t_{i1},t_{i2},\cdots,t_{ik}$ for each source $s_{i}$ with the incremental target quality, the comparison accuracy is calculated as follows:
@@ 3.3 Preliminary Study on SummEval
We conduct a preliminary study on SummEval to briefly observe the performance of LLM-as-a-Judge under different NLG meta-evaluation approaches, with detailed settings described in Appendix A.3. The results in Table 1 have shown significant ranking differences, highlighting the complexity of meta-evaluation. Although some work (Gao et al., 2024) has analyzed different correlation measures, their focus was only on the specific evaluation stability. In contrast, our proposed meta-evaluation approach focuses on different evaluation capabilities, thereby offering better interpretability. For instance, global-perspective meta-evaluation is more suitable for scenarios requiring a coarse-grained qualitative evaluation, such as selecting high-quality training data. Local-perspective meta-evaluation is more suitable for scenarios requiring a fine-grained comparative evaluation, such as labeling data pairs for preference optimization.
LLM-as-Judges $\kappa$ $Acc$ $r$ $\rho$
GPT-4o 0.225(3) 0.835(4) 0.562(4) 0.522(2)
GPT-4o mini 0.128(6) 0.815(6) 0.521(7) 0.507(4)
GPT-4 Turbo 0.342(1) 0.839(3) 0.625(1) 0.513(3)
Llama-3.1-70B 0.114(7) 0.807(7) 0.523(6) 0.491(5)
Qwen-2.5-72B 0.217(4) 0.840(2) 0.577(2) 0.490(7)
Gemma-2-27B 0.211(5) 0.842(1) 0.533(5) 0.490(6)
Phi-4-14B 0.301(2) 0.815(5) 0.564(3) 0.532(1)
Table 1: The results and rankings of common LLMs as evaluators from our proposed two meta-evaluation perspectives and under the traditional meta-evaluation with Pearson ($r$) and Spearman ($\rho$) correlations.
@@ 4.1 Evaluation Aspect Decomposition
In common NLG benchmarks, the evaluation is often further specified across different aspects like coherence. We first decompose them into more fine-grained sub-aspects for better guidance of the error injection. Previous research on active evaluation (Liu et al., 2024d; Li et al., 2025) has also shown the aspect decomposition can improve the evaluation performance of LLMs, which may play a role similar to the chain of thought (Wei et al., 2022). To achieve the sub-aspects efficiently, we prompt OpenAI o1 (OpenAI, 2024), leveraging its strong reasoning ability to first generate candidates. Then, they undergo manual selection and refinement, yielding a set of representative sub-aspects, as shown in Tables 12 and 13 in the appendix.
@@ 4.2 Global Perspective
Since human evaluation is generally regarded as the gold standard, we adhere to the original human evaluation scale to define the different coarse-grained ratings and construct corresponding new targets. Firstly, we generate candidate targets with various qualities by simultaneously injecting different numbers of errors into references using OpenAI o1. Each error is associated with a random evaluation sub-aspect described in Section 4.1. Additionally, the references are generated by GPT-4o, and we apply a heuristic algorithm to select the most diverse portions from multiple samplings, ensuring both the high quality and diversity of references.
@@
Since each candidate target $t$ must belong to a certain rating $r$, it should basically have higher quality than anchor targets of rating $r-1$ and lower quality than anchor targets of rating $r+1$, while exhibiting similar quality to anchor targets of rating $r$. Based on this principle, we compare each candidate target with anchor targets of each rating in a pairwise manner to estimate its rating:
@@
where $|A_{r}|$ represents the number of targets in anchor set $A_{r}$, and $F$ denotes the direct pairwise comparison implemented by the LLM, returning 1 if the given quality relationship holds and 0 otherwise. Furthermore, to reduce the number of comparisons and the corresponding computational cost, we use a small part of the data in advance to approximate the number of errors corresponding to each rating for more efficient target generation.
LLM SummEval Global Topical-Chat Global Overall
Coh Con Flu Rel Avg Und Nat MCtx Int UK Avg
GPT-4o 0.677 0.601 0.595 0.737 0.653 0.947 0.667 0.756 0.625 0.713 0.742 0.697 (5)
GPT-4o mini 0.619 0.615 0.584 0.781 0.650 0.967 0.685 0.680 0.586 0.677 0.719 0.684 (6)
GPT-4 Turbo 0.621 0.570 0.569 0.708 0.617 0.930 0.593 0.693 0.711 0.720 0.730 0.673 (7)
GPT-3.5 Turbo 0.551 0.556 0.513 0.712 0.583 0.753 0.583 0.444 0.683 0.613 0.615 0.599 (12)
DeepSeek-V3 0.625 0.532 0.587 0.683 0.607 0.863 0.588 0.654 0.589 0.720 0.683 0.645 (8)
Llama-3.1-70B 0.559 0.677 0.588 0.560 0.596 0.771 0.625 0.607 0.580 0.493 0.615 0.606 (11)
Qwen-2.5-72B 0.737 0.818 0.786 0.733 0.769 0.943 0.820 0.737 0.783 0.760 0.809 0.789 (1)
Gemma-2-27B 0.601 0.710 0.532 0.724 0.642 0.913 0.507 0.552 0.558 0.597 0.626 0.634 (9)
Phi-4-14B 0.649 0.608 0.667 0.659 0.646 0.903 0.644 0.694 0.783 0.753 0.756 0.701 (4)
Auto-J-13B 0.290 0.274 0.169 0.344 0.269 0.026 0.168 0.201 0.302 0.021 0.144 0.207 (16)
CRITIQUELLM-6B 0.548 0.608 0.318 0.654 0.532 0.418 0.113 0.239 0.292 0.434 0.299 0.415 (14)
Prometheus-13B 0.333 0.473 0.434 0.407 0.412 0.223 0.161 0.137 0.310 0.334 0.233 0.322 (15)
Prometheus-2-7B 0.638 0.634 0.401 0.656 0.582 0.585 0.393 0.584 0.558 0.593 0.542 0.562 (13)
Prometheus-2-8x7B 0.690 0.755 0.548 0.613 0.652 0.620 0.452 0.605 0.767 0.564 0.602 0.627 (10)
Themis-8B 0.797 0.827 0.741 0.797 0.791 0.852 0.729 0.628 0.702 0.317 0.645 0.718 (3)
CompassJudger-32B 0.782 0.863 0.705 0.840 0.797 0.863 0.643 0.649 0.869 0.577 0.720 0.759 (2)
Table 2: The results of the linear weighted $\kappa$ for different LLMs on our new global-perspective benchmarks, with the overall ranking shown in parentheses. The abbreviations represent different evaluation aspects and their averages.
@@ 4.3 Local Perspective
As for the local perspective, we aim to construct a sequence of targets for each source that follows a clear quality ranking without explicit ratings. To achieve this, we modify the previous error injection method to an iterative process, where we start from the reference and insert one error at a time while unchanging other content. Each error-injected target serves as the input for the next iteration, with each error also randomly corresponding to an evaluation sub-aspect. Since cumulative errors objectively ensure a gradual decline in quality, the sequence from the reference to the last target after the error injection exhibits an order of descending quality. Moreover, free from the constraints of the original human evaluation scale, we can customize the number of iterations to make the sequence contain more targets than the number of different ratings in the scale, thereby ensuring more fine-grained quality differences. The iterative injection process is also implemented by OpenAI o1 with elaborate instructions. Other details of our automatic benchmark construction are included in Appendix C.1.

### v2

@@ 1 Introduction
Natural Language Generation (NLG) has long been crucial in natural language processing research, encompassing common tasks such as text summarization and dialogue response generation. Unlike other tasks, such as question answering and mathematical reasoning, which have a single standard answer, making it convenient to evaluate the model output by direct matching, NLG tasks are inherently open-ended and lack a unique correct output. Therefore, they require more robust and flexible evaluation metrics. Traditional metrics primarily include string matching-based methods, such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), as well as semantic representation-based methods like BERTScore (Zhang et al., ) and BARTScore (Yuan et al., 2021). However, with the advancements of language models, these metrics have gradually been found insufficient to meet requirements. Moreover, some studies (Sai et al., 2021; He et al., 2023) have highlighted their deficiencies in robustness.
@@
Nevertheless, we argue that traditional NLG meta-evaluation used to assess the performance of metrics has certain limitations, preventing a comprehensive understanding of their evaluation capabilities. Firstly, when measuring the consistency between metrics and human evaluations, multiple human ratings for each sample are typically averaged directly. However, human ratings are not continuous values and may be inconsistent among different annotators, and the quality intervals between adjacent ratings are not uniform (Sullivan and Artino Jr, 2013). Therefore, this seemingly intuitive averaging aggregation is not such reasonable. In addition, the choice of correlation measures commonly used in traditional meta-evaluation can significantly impact the performance assessment of different metrics (Perrella et al., 2024b; Gao et al., 2024). It is unclear which correlation measure should be selected in different NLG evaluation scenarios. Furthermore, many widely used NLG evaluation benchmarks suffer from issues including outdated generation systems and potential data contamination, further undermining the reliability of meta-evaluation. More detailed discussions of these limitations are included in Section 3.
To address these issues, we propose a new dual-perspective NLG meta-evaluation framework. Specifically, the meta-evaluation from the global perspective formulates evaluation as an ordinal classification task, assessing the ability of metrics to judge the coarse-grained quality ratings of targets. On the other hand, the local-perspective meta-evaluation formulates evaluation as an adjacent pairwise comparison task, assessing the metrics in distinguishing between targets with fine-grained quality differences. These two perspectives focus on different evaluation capabilities, forming a more interpretable NLG meta-evaluation approach compared to the traditional meta-evaluation.
Furthermore, we introduce a method of automatic benchmark construction to better implement our proposed dual-perspective meta-evaluation. Our new benchmarks can be automatically constructed based on any existing NLG evaluation benchmark with new content, which avoids the cost of additional human annotations and potential data contamination. To comprehensively assess the evaluation capabilities of emerging LLM-as-a-Judge, we further experiment with 16 representative LLMs, including both the general-purpose and specifically fine-tuned ones, based on our new NLG meta-evaluation framework. The results indicate LLMs have their respective proficiency in different evaluation capabilities, demonstrating the effectiveness and necessity of our framework.
@@
    We conduct extensive experiments with 16 representative LLMs, analyzing their respective proficiency in different evaluation capabilities. Our code and data are released in https://github.com/PKU-ONELab/NLG-DualEval to facilitate related research.
@@ 3 Dual-Perspective Meta-Evaluation
Recent studies (Perrella et al., 2024b; Gao et al., 2024) have investigated the impact of different consistency measures on NLG meta-evaluation, but less attention has been paid to the aggregation method. The implementation of directly averaging ratings from multiple human evaluators has been followed by default in the previous related work. However, we have doubts about this seemingly self-evident approach from two main aspects. We take the widely-used benchmark, SummEval (Fabbri et al., 2021), as an example, where human evaluation is conducted using a 5-point Likert scale with three annotators.
@@
The additional quality levels generated by averaging aggregation are not necessarily valid. The averaged ratings with the 1-5 scale and three annotators can have up to 13 distinct values, leading to some new ratings besides the original five ratings, such as rating $\frac{4}{3}$. To verify whether these new ratings really reflect finer-grained quality levels beyond the original ratings, we conduct a human re-evaluation through pairwise comparison on SummEval. Specifically, we sample specific target pairs as the first test group, where the two averaged ratings involved in each pair are adjacent and belong to new ratings and original ratings, respectively, such as ratings $\frac{4}{3}$ and 1. As a contrast, we additionally sample the second test groups, where the two averaged ratings involved are adjacent original ratings, such as ratings 2 and 1. More experimental details are included in Appendix A.1. The proportions of pairs with consistent judgments between our re-evaluations and the original averaged human ratings from SummEval are only 42% for the first group, while 88% for the second group. The results indicate that there are indeed definite quality differences among the original ratings, whereas this is not the case if adding the new ratings generated by averaging aggregation.
@@ 3.1 Global Perspective
$\mathrm{CEM}(\{h_{ij}\}_{i=1,j=1}^{n,m},\{x_{ij}\}_{i=1,j=1}^{n,m},l)$
where $l$ represents the number of different coarse-grained ratings, and $\mathrm{CEM}$ represents the Closeness Evaluation Measure, which is an effective ordinal classification metric proposed by Amigó et al. (2020). More importantly, it does not require the assumption of equal intervals between adjacent categories, aligning with our previous discussions on the issues of the averaging aggregation method. More details and descriptions of $\mathrm{CEM}$ are included in Appendix A.2.
@@ 3.2 Local Perspective
On the other hand, in local-perspective meta-evaluation, the focus shifts to the ability to perceive fine-grained quality differences, thereby complementing the global perspective and rewarding those metrics with good discriminative ability. More specifically, given the targets with different qualities, where some of them may have the same coarse-grained rating, the concern here is whether they can be correctly ranked rather than whether they can be rated correctly. Furthermore, through a high-density target sequence for each source that contains significantly more than $l$ targets with different qualities, we can enable the more fine-grained meta-evaluation. The performance of the evaluation metric is assessed by comparing their evaluation scores for each pair of adjacent targets in the sequence sorted by target quality, which is the most challenging case. Given a target sequence $t_{i1},t_{i2},\cdots,t_{ik}$ for each source $s_{i}$ with the incremental target quality, the comparison accuracy is calculated as follows:
@@ 3.3 Preliminary Study on SummEval
We conduct a preliminary experiment on SummEval using our proposed meta-evaluation approaches, along with two traditional correlation measures, with detailed settings provided in Appendix A.3. As shown in Table 1, different meta-evaluation approaches lead to differentiated performance rankings of evaluators, highlighting the complexity of meta-evaluation and the necessity of enhancing its interpretability. Although some work (Gao et al., 2024) has analyzed different correlation measures, their focus was primarily on evaluation stability in specific aspects. In contrast, our proposed meta-evaluation approach focuses on different evaluation capabilities, thereby offering more intuitive guidance for practical evaluation scenarios. For instance, global-perspective meta-evaluation is more suitable for scenarios requiring coarse-grained qualitative judgments, such as selecting high-quality training data. Local-perspective meta-evaluation is more suitable for scenarios requiring a fine-grained comparative evaluation, such as labeling data pairs for preference optimization.
LLM-as-Judges $\mathrm{CEM}$ AccAcc $r$ $\rho$
GPT-4o 0.580(3) 0.835(4) 0.562(4) 0.522(2)
GPT-4o mini 0.393(7) 0.815(6) 0.521(7) 0.507(4)
GPT-4 Turbo 0.741(1) 0.839(3) 0.625(1) 0.513(3)
Llama-3.1-70B 0.402(6) 0.807(7) 0.523(6) 0.491(5)
Qwen-2.5-72B 0.518(5) 0.840(2) 0.577(2) 0.490(7)
Gemma-2-27B 0.527(4) 0.842(1) 0.533(5) 0.490(6)
Phi-4-14B 0.680(2) 0.815(5) 0.564(3) 0.532(1)
Table 1: The results and rankings of common LLMs as evaluators from our proposed two meta-evaluation perspectives (columns 2 and 3), or using the traditional Pearson ($r$) and Spearman ($\rho$) correlation coefficients (columns 4 and 5).
Evaluation Aspect Decomposed Sub-aspects
Table 2: An example of the evaluation aspect coherence and its decomposed fine-grained sub-aspects in SummEval. The complete information is presented in Tables 13 and 14 in the appendix.
@@ 4.1 Evaluation Aspect Decomposition
In common NLG benchmarks, the evaluation is often further specified across different evaluation aspects like coherence. We first decompose them into more fine-grained sub-aspects for better guidance of the error injection. Previous research on active evaluation (Liu et al., 2024d; Li et al., 2025) has also shown the aspect decomposition can improve the evaluation performance of LLMs, which may play a role similar to the chain of thought (Wei et al., 2022). To achieve the sub-aspects efficiently, we first prompt OpenAI o1 (OpenAI, 2024), leveraging its strong reasoning ability to generate candidate sub-aspects. Then, they undergo manual selection and refinement, yielding a set of representative sub-aspects, as exemplified in Table 2.
LLM SummEval Global Topical-Chat Global Overall
Coh Con Flu Rel Avg Und Nat MCtx Int UK Avg
GPT-4o 0.765 0.708 0.702 0.801 0.744 0.978 0.815 0.875 0.792 0.888 0.870 0.807 (5)
GPT-4o mini 0.738 0.726 0.712 0.835 0.753 0.985 0.833 0.826 0.767 0.877 0.858 0.805 (6)
GPT-4 Turbo 0.731 0.691 0.690 0.785 0.724 0.972 0.774 0.841 0.849 0.889 0.865 0.795 (7)
GPT-3.5 Turbo 0.673 0.667 0.653 0.779 0.693 0.905 0.772 0.699 0.829 0.844 0.810 0.751 (12)
DeepSeek-V3 0.739 0.665 0.710 0.763 0.719 0.947 0.769 0.819 0.771 0.894 0.840 0.779 (10)
Llama-3.1-70B 0.700 0.779 0.712 0.693 0.721 0.914 0.800 0.789 0.766 0.810 0.816 0.768 (11)
Qwen-2.5-72B 0.814 0.868 0.838 0.800 0.830 0.978 0.910 0.862 0.886 0.904 0.908 0.869 (1)
Gemma-2-27B 0.738 0.787 0.698 0.790 0.753 0.966 0.724 0.750 0.751 0.842 0.806 0.780 (9)
Phi-4-14B 0.761 0.721 0.775 0.752 0.752 0.963 0.823 0.837 0.887 0.901 0.882 0.817 (4)
Auto-J-13B 0.615 0.610 0.569 0.644 0.610 0.614 0.602 0.589 0.666 0.612 0.617 0.613 (16)
CRITIQUELLM-6B 0.710 0.737 0.615 0.765 0.707 0.769 0.577 0.602 0.670 0.776 0.679 0.693 (14)
Prometheus-13B 0.576 0.653 0.645 0.613 0.622 0.692 0.569 0.557 0.638 0.736 0.639 0.630 (15)
Prometheus-2-7B 0.739 0.733 0.639 0.746 0.714 0.836 0.689 0.774 0.751 0.839 0.778 0.746 (13)
Prometheus-2-8x7B 0.784 0.828 0.714 0.724 0.762 0.849 0.740 0.793 0.878 0.827 0.818 0.790 (8)
Themis-8B 0.852 0.873 0.809 0.847 0.845 0.941 0.859 0.805 0.843 0.729 0.835 0.840 (3)
CompassJudger-32B 0.844 0.899 0.797 0.881 0.855 0.946 0.818 0.819 0.928 0.835 0.869 0.862 (2)
Table 3: The results of the $\mathrm{CEM}$ metric for different LLMs on our new global-perspective benchmarks, with the overall performance ranking shown in parentheses. The abbreviations in columns 2-12 represent different evaluation aspects as shown in Tables 13 and 14, and the average results of each benchmark.
@@ 4.2 Global Perspective
Since human evaluation is generally regarded as the gold standard, we adhere to the original human evaluation scale to define the different coarse-grained ratings and construct corresponding new targets. Firstly, we generate candidate targets with various qualities by simultaneously injecting different numbers of errors into references using OpenAI o1 and the prompt in Table 18. Each error is associated with a random evaluation sub-aspect. Additionally, to ensure both the high quality and diversity of the references, they are generated by GPT-4o with multiple samplings using the prompt in Table 12. We apply a heuristic algorithm to select the most diverse portions based on the similarity measure.
@@
Since each candidate target $t$ must belong to a certain rating $r$, it should basically have higher quality than anchor targets of rating r1r-1 and lower quality than anchor targets of rating $r+1$, while exhibiting similar quality to anchor targets of rating $r$. Based on this principle, we compare each candidate target with anchor targets of each rating in a pairwise manner to estimate its rating:
@@
where $|A_{r}|$ represents the number of targets in anchor set $A_{r}$, and $F$ denotes the direct pairwise comparison implemented by the LLM, returning 1 if the given quality relationship holds and 0 otherwise. Furthermore, to reduce the number of comparisons and the corresponding computational cost, we use a small part of the data to approximate the error number that corresponds to each rating to guide the more efficient target generation. More details of the anchoring are included in Appendix C.1.
@@ 4.3 Local Perspective
As for the local perspective, we aim to construct a sequence of targets for each source that follows a clear quality ranking without explicit ratings. To achieve this, we modify the previous error injection method to an iterative process, where we start from the reference and insert one error at a time while unchanging other content. Each error-injected target serves as the input for the next iteration, with each error also randomly corresponding to an evaluation sub-aspect. Since cumulative errors objectively ensure a gradual decline in quality, the sequence from the reference to the last target after the error injection exhibits an order of descending quality. Moreover, free from the constraints of the original human evaluation scale, we can customize the number of iterations to make the sequence contain more targets than the number of different ratings in the scale, thereby ensuring more fine-grained quality differences. The iterative injection process is also implemented by OpenAI o1 with elaborate instructions in Table 19. Other details of our benchmark construction are provided in Appendix C.1.